OpenStudy (anonymous):

What is the equation of the line that passes through the points (-2, 1) and (1, 10)?

OpenStudy (anonymous):

use two point formula

OpenStudy (ghazi):

(y-1)=(10-1)/(1+2)*(x+1) y-1= 9/3*(x-1)

OpenStudy (anonymous):

yes the same

OpenStudy (ghazi):

y-1=3*(x-1) y-1=3x-3 rest you can do :)

OpenStudy (anonymous):

:) thank you

OpenStudy (ghazi):

:)